# NASTaR: NovaSAR Automated Ship Target Recognition Dataset

### v1

@@ Abstract
Synthetic Aperture Radar (SAR) offers a unique capability for all-weather, space-based maritime activity monitoring by capturing and imaging strong reflections from ships at sea. A well-defined challenge in this domain is ship type classification. Due to the high diversity and complexity of ship types, accurate recognition is difficult and typically requires specialized deep learning models. These models, however, depend on large, high-quality ground-truth datasets to achieve robust performance and generalization. Furthermore, the growing variety of SAR satellites operating at different frequencies and spatial resolutions has amplified the need for more annotated datasets to enhance model accuracy. To address this, we present the NovaSAR Automated Ship Target Recognition (NASTaR) dataset. This dataset comprises of 3415 ship patches extracted from NovaSAR S-band imagery, with labels matched to AIS data. It includes distinctive features such as 23 unique classes, inshore/offshore separation, and an auxiliary wake dataset for patches where ship wakes are visible. We validated the dataset’s applicability across prominent ship-type classification scenarios using benchmark deep learning models. Results demonstrate over 60% accuracy for classifying four major ship types, over 70% for a three-class scenario, more than 75% for distinguishing cargo from tanker ships, and over 87% for identifying fishing vessels. The NASTaR dataset is available at https://10.5523/bris, while relevant codes for benchmarking and analysis are available at https://github.com/benyaminhosseiny/nastar.
@@ I Introduction
Synthetic Aperture Radar (SAR) is a coherent imaging technique capable of high-resolution imaging regardless of weather or lighting conditions. Unlike optical sensors, SAR ability to operate in day-and-night scenarios and penetrate cloud cover makes it indispensable for maritime monitoring tasks such as ship detection, classification, and tracking [SARManual2004].
SAR systems operate across various frequency bands, such as X, C, L, and more recently, the S-band, each offering distinct trade-offs. Higher-frequency bands (e.g., X and C bands) provide finer spatial resolution, enabling detailed object delineation but suffer from limited penetration and lower SNR in cluttered environments. Lower-frequency bands (e.g., the L-band) offer deeper penetration and broader swath coverage, but at the cost of reduced spatial resolution. S-band SAR, exemplified by platforms like NovaSAR and NiSAR, offers a compromise between resolution and coverage. Its moderate wavelength allows for improved SNR and wider swath acquisition compared to the X- and C bands, while retaining sufficient resolution for object-level analysis [Kamirul2024]. These characteristics make S-band SAR particularly promising for large-scale maritime surveillance, where both spatial detail and temporal coverage are critical.
SAR imagery is fundamentally different from optical imagery due to its coherent nature and side-looking geometry. Key phenomena such as speckle noise, layover, foreshortening, and shadowing complicate interpretation. In addition, the backscattering response of maritime targets is influenced by factors such as hull geometry, surface roughness, and sea state, leading to highly variable target signatures. These complexities pose significant challenges for automatic target recognition (ATR) in SAR imagery. Traditional approaches based on physical scattering models or handcrafted features often fail to generalize across varying acquisition conditions and target types [awais2025survey].
@@
Benchmark datasets play a critical role in advancing DL research by providing standardized evaluation protocols and facilitating reproducibility. While several SAR datasets exist for ship detection and classification, most are based on X- or C-band imagery and focus on binary detection or coarse classification tasks [zhang2024development]. There is a notable gap in publicly available datasets tailored to S-band SAR imagery, especially for fine-grained ship-type classification. Given the unique imaging characteristics of the S-band and its growing adoption in maritime surveillance, a dedicated benchmark is essential to support model development, cross-band generalization studies, and operational deployment.
@@ II-A Existing SAR Ship Recognition Datasets
Publicly available datasets for ship ATR using SAR imagery remain limited. One of the earliest is the HR-SAR dataset [xing2013ship], released in 2013, comprising 450 high-resolution (2×1.5 m, X-band) TerraSAR-X images across three ship categories. Another widely used resource is OpenSARShip [huang2017opensarship, li2017opensarship2]. Derived from Sentinel-1 data in SLC and GRD formats, under VH and VV polarizations A key limitation of this dataset is significant class imbalance, which hampers classification robustness and generalization. xView3-SAR [paolo2022xview3] is another dataset focused on detecting dark vessels, making it suitable for identifying illegal activities using Sentinel-1 GRD imagery. Other notable datasets originate from the Gaofen satellite series, known for their high spatial resolution and detailed backscattering. Examples include FUSAR [hou2020fusar] and MTCD [ma2018ship]. Note that none of the above datasets include S-band SAR imagery.
Although ship wake patterns are useful for identifying smaller vessels that are difficult to detect in low-resolution SAR imagery, datasets dedicated to wake detection and identification remain extremely limited. Most existing datasets concentrate on wake detection, rather than classification. For instance, OpenSARWake [xu2024opensarwake] supports wake detection tasks and includes samples from Sentinel-1, Gaofen, and ALOS-PALSAR satellites. However, to the best of our knowledge, there is no publicly available dataset specifically designed for ship identification based on wake patterns. The only notable exception is SynthWakeSAR [rizaev2022synthwakesar], which simulates SAR images of ship wakes based on hydrological modeling of the sea surface.
@@ II-B Deep learning Techniques in Ship Target Recognition
Recent research efforts have been undertaken to improve SAR ship ATR by addressing main challenges, such as class imbalance, noisy, varied shapes, and sparse backscattering. For example, to mitigate the class imbalance problem in the FUSAR and OpenSARShip datasets, Zhang et al., [zhang2021imbalanced] proposed a lightweight customized CNN model and a training process based on identifying the center of each class in the deep feature space, followed by gradually balanced sampling. To address the large shape and size variations among different ship types and the significant inter-class overlap, a multiscale feature attention mechanism was introduced in [wang2023sar]. This approach improved classification accuracy through an adaptive weighting technique that effectively emphasizes features from different scales. To enhance model robustness and reliability in ship recognition, Zheng et al., [zheng2022metaboost] developed a MetaBoost-based ensemble learning model, which demonstrated improved performance across diverse SAR scenes.
@@ III The NASTaR Dataset
Coupling the image with ground truth AIS data is a key element of NASTaR. The NovaSAR satellite carries its own AIS instrument, meant to provide complementary ship tracking data for maritime monitoring applications. Here, however, we have opted to correlate the ships in the SAR images with AIS data provided by the Danish Maritime Authority [DMA_AIS_Data] as their ground-based AIS transponders covering our area of interest are more accurate and capable of reliably capturing signals from dense ship populations. The AIS samples were filtered to the same geographical zone as the corresponding SAR scene and within a 5-minute time window around the acquisition time. In addition to the ship’s location and timestamp, the AIS dataset includes information such as speed over ground (SOG), heading angle, and vessel type and dimensions.
Each ship patch sample is generated by mapping the latitude and longitude from the AIS data to the corresponding location in the NovaSAR image. The patches are sized at 512×512 pixels, equivalent to 1280×1280 m² on the ground. Initially, 3415 ship patches were extracted. To classify these patches as either inshore or offshore, the distance from the ship’s location to the nearest landmass was calculated using Natural Earth maps [Natural_Earth_vec]. A threshold of 500 meters was used to distinguish between inshore and offshore patches, resulting in 1960 inshore and 1455 offshore patches. The calculated distance to shore per sample is provided so as to enable users to define custom inshore/offshore splits as needed.
@@ IV Benchmark Experiments
Fishing vessels are used to harvest seafood and are essential to the global food supply chain.Monitoring and detecting fishing vessels is especially important due to the prevalence of illegal, unreported, and unregulated (IUU) fishing, which threatens marine biodiversity and the sustainability of global fish stocks, while costing legitimate fishermen and governments billions of dollars [paolo2024satellite]. Detecting these vessels is challenging, particularly when they operate in remote areas or disable tracking systems. Their relatively small size and erratic movement patterns further complicate identification using satellite SAR imagery. Passenger ships, including ferries and cruise liners, are designed to transport people and serve as key components of both public transportation and the tourism sector. Monitoring these ship types is essential for ensuring operational efficiency, environmental compliance, maritime safety, and the protection of marine ecosystems.
We considered five classification scenarios derived from combinations of four primary ship types using benchmark deep learning models, including ResNet, and ResNeXt variants[He2015, Xie2016], DenseNet121 [huang2017densely], EfficientNet [tan2021efficientnetv2], and Vision Transformers [tiny_vit]. For brevity, only average results are reported . The framework, implemented in PyTorch, used Adam optimization with weighted cross-entropy loss to address class imbalance. Training employed a batch size of 16, learning rates between 0.0001–0.0005, and linear decay after five epochs without validation improvement. Ten-fold cross-validation ensured robustness, with 70% of data for training per fold.

### v2

# NASTaR: A NovaSAR-Based Automated Ship Target Recognition Dataset

@@ Abstract
Synthetic Aperture Radar (SAR) offers a unique capability for all-weather, space-based maritime activity monitoring by capturing and imaging strong reflections from ships at sea. A well-defined challenge in this domain is ship type classification. Due to the high diversity and complexity of ship types, accurate recognition is difficult and typically requires specialized deep learning models. These models, however, depend on large, high-quality ground-truth datasets to achieve robust performance and generalization. Furthermore, the growing variety of SAR satellites operating at different frequencies and spatial resolutions has amplified the need for more annotated datasets to enhance model accuracy. To address this, we present the NovaSAR Automated Ship Target Recognition (NASTaR) dataset. This dataset comprises of 3415 ship patches extracted from NovaSAR S-band imagery, with labels matched to AIS data. It includes distinctive features such as 23 unique classes, inshore/offshore separation, and an auxiliary wake dataset for patches where ship wakes are visible. We validated the dataset’s applicability across prominent ship-type classification scenarios using benchmark deep learning models. Results demonstrate over 60% accuracy for classifying four major ship types, over 70% for a three-class scenario, more than 75% for distinguishing cargo from tanker ships, and over 87% for identifying fishing vessels. The NASTaR dataset is available at https://doi.org/10.5523/bris.2tfa6x37oerz2lyiw6hp47058, while relevant codes for benchmarking and analysis are available at https://github.com/benyaminhosseiny/nastar.
@@ I Introduction
Synthetic Aperture Radar (SAR) is a coherent imaging technique capable of high-resolution imaging regardless of weather or lighting conditions. Unlike optical sensors, SAR ability to operate in day-and-night scenarios and penetrate cloud cover makes it indispensable for maritime monitoring tasks such as ship detection, classification, and tracking [1].
SAR systems operate across various frequency bands, such as X, C, L, and more recently, the S-band, each offering distinct trade-offs. Higher-frequency bands (e.g., X and C bands) provide finer spatial resolution, enabling detailed object delineation but suffer from limited penetration and lower SNR in cluttered environments. Lower-frequency bands (e.g., the L-band) offer deeper penetration and broader swath coverage, but at the cost of reduced spatial resolution. S-band SAR, exemplified by platforms like NovaSAR and NiSAR, offers a compromise between resolution and coverage. Its moderate wavelength allows for improved SNR and wider swath acquisition compared to the X- and C bands, while retaining sufficient resolution for object-level analysis [2]. These characteristics make S-band SAR particularly promising for large-scale maritime surveillance, where both spatial detail and temporal coverage are critical.
SAR imagery is fundamentally different from optical imagery due to its coherent nature and side-looking geometry. Key phenomena such as speckle noise, layover, foreshortening, and shadowing complicate interpretation. In addition, the backscattering response of maritime targets is influenced by factors such as hull geometry, surface roughness, and sea state, leading to highly variable target signatures. These complexities pose significant challenges for automatic target recognition (ATR) in SAR imagery. Traditional approaches based on physical scattering models or handcrafted features often fail to generalize across varying acquisition conditions and target types [3].
@@
Benchmark datasets play a critical role in advancing DL research by providing standardized evaluation protocols and facilitating reproducibility. While several SAR datasets exist for ship detection and classification, most are based on X- or C-band imagery and focus on binary detection or coarse classification tasks [4]. There is a notable gap in publicly available datasets tailored to S-band SAR imagery, especially for fine-grained ship-type classification. Given the unique imaging characteristics of the S-band and its growing adoption in maritime surveillance, a dedicated benchmark is essential to support model development, cross-band generalization studies, and operational deployment.
@@ II-A Existing SAR Ship Recognition Datasets
Publicly available datasets for ship ATR using SAR imagery remain limited. One of the earliest is the HR-SAR dataset [5], released in 2013, comprising 450 high-resolution (2×1.5 m, X-band) TerraSAR-X images across three ship categories. Another widely used resource is OpenSARShip [6, 7]. Derived from Sentinel-1 data in SLC and GRD formats, under VH and VV polarizations A key limitation of this dataset is significant class imbalance, which hampers classification robustness and generalization. xView3-SAR [8] is another dataset focused on detecting dark vessels, making it suitable for identifying illegal activities using Sentinel-1 GRD imagery. Other notable datasets originate from the Gaofen satellite series, known for their high spatial resolution and detailed backscattering. Examples include FUSAR [9] and MTCD [10]. Note that none of the above datasets include S-band SAR imagery.
Although ship wake patterns are useful for identifying smaller vessels that are difficult to detect in low-resolution SAR imagery, datasets dedicated to wake detection and identification remain extremely limited. Most existing datasets concentrate on wake detection, rather than classification. For instance, OpenSARWake [11] supports wake detection tasks and includes samples from Sentinel-1, Gaofen, and ALOS-PALSAR satellites. However, to the best of our knowledge, there is no publicly available dataset specifically designed for ship identification based on wake patterns. The only notable exception is SynthWakeSAR [12], which simulates SAR images of ship wakes based on hydrological modeling of the sea surface.
@@ II-B Deep learning Techniques in Ship Target Recognition
Recent research efforts have been undertaken to improve SAR ship ATR by addressing main challenges, such as class imbalance, noisy, varied shapes, and sparse backscattering. For example, to mitigate the class imbalance problem in the FUSAR and OpenSARShip datasets, Zhang et al., [13] proposed a lightweight customized CNN model and a training process based on identifying the center of each class in the deep feature space, followed by gradually balanced sampling. To address the large shape and size variations among different ship types and the significant inter-class overlap, a multiscale feature attention mechanism was introduced in [14]. This approach improved classification accuracy through an adaptive weighting technique that effectively emphasizes features from different scales. To enhance model robustness and reliability in ship recognition, Zheng et al., [15] developed a MetaBoost-based ensemble learning model, which demonstrated improved performance across diverse SAR scenes.
@@ III The NASTaR Dataset
Coupling the image with ground truth AIS data is a key element of NASTaR. The NovaSAR satellite carries its own AIS instrument, meant to provide complementary ship tracking data for maritime monitoring applications. Here, however, we have opted to correlate the ships in the SAR images with AIS data provided by the Danish Maritime Authority [16] as their ground-based AIS transponders covering our area of interest are more accurate and capable of reliably capturing signals from dense ship populations. The AIS samples were filtered to the same geographical zone as the corresponding SAR scene and within a 5-minute time window around the acquisition time. In addition to the ship’s location and timestamp, the AIS dataset includes information such as speed over ground (SOG), heading angle, and vessel type and dimensions.
Each ship patch sample is generated by mapping the latitude and longitude from the AIS data to the corresponding location in the NovaSAR image. The patches are sized at 512×512 pixels, equivalent to 1280×1280 m² on the ground. Initially, 3415 ship patches were extracted. To classify these patches as either inshore or offshore, the distance from the ship’s location to the nearest landmass was calculated using Natural Earth maps [17]. A threshold of 500 meters was used to distinguish between inshore and offshore patches, resulting in 1960 inshore and 1455 offshore patches. The calculated distance to shore per sample is provided so as to enable users to define custom inshore/offshore splits as needed.
@@ IV Benchmark Experiments
Fishing vessels are used to harvest seafood and are essential to the global food supply chain.Monitoring and detecting fishing vessels is especially important due to the prevalence of illegal, unreported, and unregulated (IUU) fishing, which threatens marine biodiversity and the sustainability of global fish stocks, while costing legitimate fishermen and governments billions of dollars [18]. Detecting these vessels is challenging, particularly when they operate in remote areas or disable tracking systems. Their relatively small size and erratic movement patterns further complicate identification using satellite SAR imagery. Passenger ships, including ferries and cruise liners, are designed to transport people and serve as key components of both public transportation and the tourism sector. Monitoring these ship types is essential for ensuring operational efficiency, environmental compliance, maritime safety, and the protection of marine ecosystems.
We considered five classification scenarios derived from combinations of four primary ship types using benchmark deep learning models, including ResNet, and ResNeXt variants[19, 20], DenseNet121 [21], EfficientNet [22], and Vision Transformers [23]. For brevity, only average results are reported. The framework, implemented in PyTorch, used Adam optimization with weighted cross-entropy loss to address class imbalance. Training employed a batch size of 16, learning rates between 0.0001–0.0005, and linear decay after five epochs without validation improvement. Ten-fold cross-validation ensured robustness, with 70% of data for training per fold.